OpenStudy (anonymous):

4x cos x (a) lim x→0 f(x) = (b) lim x→π/3 f(x

OpenStudy (anonymous):

do you know much about limits @shorouq91

OpenStudy (anonymous):

i gont part a but i dont know who to do part b

OpenStudy (anonymous):

ok, what was your answer to part a, just to be sure

OpenStudy (anonymous):

0

OpenStudy (anonymous):

part b 4*pi/3 cos(1/2)

OpenStudy (anonymous):

u were almost there. AS cos x approches \[\frac{ \pi}{ 3 }\], it becomes half. Not \[\cos \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

so, u can write the solution as\[4\frac{ \pi }{ 3} (\frac{ 1 }{ 2 }= \frac{ 2\pi }{ 3 }\]

OpenStudy (anonymous):

i dont under stand it

OpenStudy (anonymous):

sorry i forgot to close the parenthesis . ok, let's start\[4xcosx =(\lim_{x \rightarrow \frac{ \pi }{ 3}}4x) ( \lim_{x \rightarrow \frac{ \pi }{ 3}}cosx)\]

OpenStudy (anonymous):

do you understand that part

OpenStudy (anonymous):

@shorouq91

OpenStudy (anonymous):

not much

OpenStudy (anonymous):

the limit of 4x multiplied by the limit of cosx, we will deal with the limits in two parts, 4x, and cosx, and instead of x, we plug in pi/3 like u did, such that u will have\[4\frac{ \pi }{ 3}\] on the first part. on the second part - QUESTION IS, WHAT IS PI/3 IN DEGREES, DO U KNOW?

OpenStudy (anonymous):

@shorouq91 , what z pi/3?

OpenStudy (anonymous):

in degrees?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

pi = 180 degrees, divide that by three, u will get 60 degrees, so, as x approaches 60 degrees, what what is cos 60?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

so, the second part which is cosx, as x approaches pi/3, it becomes half!, now we have:\[(4\frac{ \pi }{ 3})(\frac{ 1 }{ 2 }) = \frac{ 2\pi }{ 3 },\] which is the solution to your second part!

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

im stock in the last part

OpenStudy (anonymous):

if we are to find the limit, as x approaches pi/3, of a function like the one above, that has two parts, 4x and cosx, we plug in pi/3 in seperately, and multiply the limits when we find them, so the limit for 4x was, as x approaches pi/3= 4(pi/3) [ instead of x, we put in the limit. and the second part, also, as x approaches pi/3, which is 60 degrees, it becomes cos 60=1/2 then 4(pi/3) x 1/2 [remember it was 4x times cosx!

OpenStudy (anonymous):

4(pi/3) x 1/2 = 2pi/3

OpenStudy (anonymous):

oky got it thank you so muchh i took to much time

OpenStudy (anonymous):

always a pleasure. But did you get the trick here? You plug in the limit, but then, u must know that cos pi/3 = 1/2!